OpenStudy (anonymous):

What is the slope between the two points (8, 15) and (3, 5)?

OpenStudy (allank):

Hint: slope (gradient) = (x1-x2)/(y1-y2) where (x1,y1) and (x2,y2) are the two points

OpenStudy (anonymous):

i dont want a hint i want an answer please :)

OpenStudy (anonymous):

I'm sorry, but this is a tutoring site. Not a get answers site. We have to guide you to your answer.

OpenStudy (allank):

@Calcmathlete has a point =)

OpenStudy (anonymous):

well, can you please cut me a little slack

OpenStudy (tiffanymak1996):

15-5/8-3 = 10/5 = 2

OpenStudy (anonymous):

thank you so much tiffanymak1996

OpenStudy (tiffanymak1996):

just learn the method, or else in exams you won't be able to answer. y2-y1/ x2-x1, it's not that hard just minus the points.